# Residual Diffusion Bridge Model for Image Restoration

### v1

@@ 1 Introduction
Universal image restoration emerges as a unified paradigm integrating the perception, representation, and elimination of diverse degradations [18, 64], with the aim of restoring high-quality images from the degraded low-quality observation. It typically encompasses a broad range of classical tasks, including denoising, deraining, dehazing, super-resolving, and others [7, 62, 24, 69, 13]. Owing to the high fidelity of restored details, it has been widely adopted as a precursor across various downstream tasks [46, 80].
@@
Diffusion models have achieved remarkable advances in universal image restoration [42]. Early methods [60, 70, 86] follow the standard diffusion paradigm [15, 22] that maps images to a Gaussian distribution, and initializes the reverse inference from pure noise. Some approaches leverage generative priors pretrained from large models [39, 45] as conditional guidance for denoising networks. Others treat various restoration tasks as inverse problems by assuming access to degradation kernels [9, 8, 63, 76]. However, the randomness of noise and reliance on specific priors compromise both stability and universality. Subsequent studies [40, 36] incorporate mean-reverting dynamics into diffusion stochastic differential equations (SDEs), clustering forward terminal states around degraded observations to retain task-relevant cues. Additionally, diffusion bridges [82] directly model point-to-point stochastic transitions between paired distributions, thereby strengthening data associations and improving restoration fidelity. Despite these advances, existing methods still rely on global noise perturbation to construct probabilistic trajectories, requiring rigid reverse denoising processes, as shown in Fig. 1. However, they fail to distinguish regions with varying degradation levels and imperfectly reconstruct intact regions, limiting restoration performance and adaptivity. Besides, a systematic and theoretical framework is absent to elucidate the intricate interconnections among existing diffusion bridge formulations.
In this work, we propose a scalable and unified diffusion bridge framework for universal image restoration, termed Residual Diffusion Bridge Model (RDBM), and conduct a comprehensive analysis of optimal distribution transitions. Specifically, we integrate the mean-reverting property of the Ornstein–Uhlenbeck SDEs with Doob’s $h$-transform [57] to guide the terminal forward states toward degraded image distribution. Meanwhile, we use residuals from given distributions to dynamically modulate the probabilistic trajectories, thereby allowing the model to learn adaptive restoration of regions with varying degradation levels while mitigating redundant reconstruction in intact areas. Moreover, we theoretically demonstrate that our formulation yields the smooth distributional transition with respect to the residual-to-noise ratio. Building upon this, we uncover the mathematical essence of mainstream diffusion bridge formulations, all of which are special cases within our framework in specific configurations. Extensive experiments are conducted to verify the superiority of our method across diverse tasks including image restoration, translation, and inpainting both qualitatively and quantitatively. Our main contributions are summarized as follows:
@@ 2 Related Work
Denoising diffusion models [61, 60] were initially developed for image generation. Methods such as DiffIR [70], DvSR [68], and SR3 [74] directly repurpose diffusion models conditioned on degraded images for image restoration task, which suffers from performance bottlenecks due to the task incompatibility. I2SB [34] and ColdDiffusion [5] bypass explicit noise perturbations and instead learn a degraded diffusion process directly via the network. Furthermore, RDDM [36], ReShift [73], and DiffUIR [81] incorporate degradation distributions into the perturbation kernels to explicitly characterize degradation-aware diffusion processes. Moreover, IRSDE [40] employs a mean-reverting process to enforce diffusion trajectories that regress toward noisy degraded images with stationary variance. Additionally, DDBM [82], BBDM [29] and GOUB [72] further apply Doob’s $h$-transform to remove terminal noise, offering a tractable alternative to pave the probability path that connects degraded and clean images, thereby achieving remarkable restoration performance. Flow matching [32, 37] discards the stochastic noise and constructs the deterministic distribution transition path, thereby facilitating the optimal transport [85]. In contrast, our RDBM introduces the residual to modulate perturbation, enabling spatially adaptive restoration. Besides, our model can extend to these diffusion bridge models and flow matching in specific settings.
@@ 3.1 Diffusion Bridge Models
Diffusion SDEs [59, 15] with drift term $\mathbf{f}(\cdot,t)$ and diffusion term $g(t)$ can be generally formulated as [23]:
@@
where $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$ is score function. Furthermore, a diffusion process defined in Eq. (1) can be driven to arrive at a particular point of interest $\boldsymbol{\mu}$ via Doob’s $h$-transform [55]:
@@ 3.2 Ornstein Uhlenbeck Process
Figure 3: Overview of mainstream diffusion bridges via SDEs, all of which are special cases of our framework. (a) OU process maps the data distribution to the prior distribution with noise. (b) OU bridge constructs probabilistic transitions between given distributions, with stochastic paths under a general noise schedule. (c) Brownian bridge constructs linear expectations of intermediate states. (d) Our RDBM leverages residuals to refine the transition with a smooth residual-to-noise ratio.
@@ 4.1 Generalized Forward Process
which is governed by two terms. The first term depends on the residual component and fixed ratio $\lambda$. The second term is entirely determined by $\theta_{t}$ series and exhibits a monotonic decline, which diverges to infinity as time $t\!\rightarrow\!0$ and converges to an infinitesimal value as time $t\!\rightarrow\!T$. Previous works [72, 82, 40] typically set $\boldsymbol{\pi}=1$, thereby performing the global noise perturbation to uniformly disrupt the overall structure of images. This induces two ill-posed issues: (i) the intact regions are redundant and imperfect reconstruction due to inevitable cumulative error in reverse process. (ii) pixel-wise numerator $x_{0}(i,j)\!-\!\boldsymbol{\mu}(i,j)]^{2}$ may exhibit discontinuous jumps, potentially distorting the smooth monotonic decay of the residual-to-noise ratio. Therefore, to maintain the dynamic equilibrium in transmission trajectories, we fix $\boldsymbol{\pi}=\mathbf{x}_{0}-\boldsymbol{\mu}$, thereby deriving our specific formulation within this framework with adaptive noise perturbation and pixel-independent residual-to-noise ratio:
@@ 4.2 Reverse Process and Training Objective
Assuming $p_{{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\boldsymbol{\mu})$ follows a Gaussian distribution centered at $m_{\dot{\theta}}(\mathbf{x}_{t},\mathbf{x}_{0})$ with a constant variance, minimizing the Kullback-Leibler divergence [25] $D_{KL}$ is equivalent to reducing the distance between the means (see Suppl. G):
@@ 4.3 Analysis
We redefine a general mean-reverting process in Eq. (9) and employ Doob’s $h$ transform to derive the generalized diffusion bridge in Eq. (10) that exhibits the property of mean-arrival. Visualization comparisons of probability paths with several diffusion processes are illustrated in Fig. 3. We configure $\boldsymbol{\pi}$ to serve as the residual component for adaptive noise perturbation, yielding a smoothly decaying residual-to-noise ratio highly compatible with image restoration. Additionally, other mainstream bridge models can be concluded in our framework, such as standard diffusion bridge [82], Brownian Bridge [29, 34], OU Bridge [72], Flow Matching [32, 37] and others, as summarized in Tab. 1. Please see Suppl. F for detailed derivations.
@@
$\theta_{t}\rightarrow 0$ $\lambda\rightarrow 0$ $\boldsymbol{\pi}=0$ Flow Matching [32, 37]
$\theta_{t}\rightarrow 0$ $\lambda$ $\boldsymbol{\pi}=1$ VE Bridge [82]
$\theta_{t}\rightarrow 0$ $\lambda\rightarrow\infty$ $\boldsymbol{\pi}=1$ VP Bridge [82]
$\theta_{t}\rightarrow 0$ $\lambda\rightarrow\frac{1}{2}$ $\boldsymbol{\pi}=1$ Brownian Bridge [29, 34]
$\theta_{t}$ $\lambda$ $\boldsymbol{\pi}=1$ OU Bridge [72]
@@ 7  Take the gradient descent step on
8 $\phantom{******}\nabla_{\theta}\|\boldsymbol{\pi}\epsilon-\boldsymbol{\pi}_{\epsilon}^{\theta}(\mathbf{x}_{t},t,\boldsymbol{\mu})\|_{1}$
@@ 5.1 Datasets and Evaluation Metrics
Image deraining. We train our model on the merged datasets from Rain13K [19] and DeRaindrop [51], which cover diverse rain streaks and densities. Evaluation is conducted on both rain- and raindrop-removal tasks using the mixed datasets [71, 77, 51, 31]. In addition, we assess zero-shot generalization on real-world datasets, including GT-Rain [4] without ground-truth for reference.
Low-light enhancement. We combine the LOL [67] and VE-LOL-L [35] datasets, which furnish real and synthetic paired samples across diverse scenes with varying illumination and noise levels. Additionally, we employ the NPE [65], MEF [44] and DICM [26] datasets to conduct zero-shot generalization on real-world scenarios.
Image desnowing. We adopt the CSD [6] dataset as the primary benchmark and evaluate real-world performance on Snow100K-Real [38], which has no ground-truth.
Image dehazing. We adopt ITS_v2 [27] and D-HAZY [10] as training benchmarks, encompassing diverse scenes under varying haze densities. The outdoor subset SOTS [27] is used for evaluation, while real-world generalization is assessed on Dense-Haze [2], NHRW [78], and NH-HAZE [3].
Image deblurring. We use the GoPro [48] dataset to perform deblurring tasks, which contains various levels of blur obtained by averaging the clear images captured in very short intervals. To further validate the generalizability, we perform zero-shot restoration on the RealBlur [53] dataset.
Benchmarks are evaluated using peak signal-to-noise ratio (PSNR) [16], structural similarity (SSIM) [66], natural image quality evaluator (NIQE) [26] in RGB space, and the learned perceptual image patch similarity (LPIPS) [58] in feature space. For fairness, we compare our method with several universal restoration methods [75, 28, 50, 79, 40, 20, 41, 72, 11, 12, 52, 30, 43], which are all re-implemented on the mixed datasets for comparisons.
@@ 5.2 Implementation Details
Our method is trained using 8 Nvidia A800 GPUs with PyTorch [49] framework for 128h. Adam optimizer and L1 loss are employed for 500k iterations with a learning rate of $1\text{e}\!-\!4$. We set the batch size as $20$ and distribute it evenly to each task. We randomly crop patches of size $256\times 256$ from the original image as network input for training and use 10 timesteps for full-resolution testing. We utilize U-Net  [56] architecture as network backbone. We change the channel number of the hidden layers $C$ to obtain different versions with varied parameter quantities:
@@ 5    $\mathbf{x}_{T}=\boldsymbol{\mu}$
Restomer [75] 2022 28.54 0.847 21.75 0.742 28.53 0.919 26.54 0.924 26.44 0.799 27.61 0.869 26.09 140.99
AirNet [28] 2022 24.78 0.774 13.05 0.485 25.80 0.885 18.53 0.827 25.76 0.782 24.01 0.809 5.76 301.27
Prompt-IR [50] 2023 28.97 0.856 20.97 0.733 29.52 0.938 25.80 0.929 26.25 0.797 27.89 0.878 32.96 158.14
ProRes [43] 2023 22.42 0.752 20.31 0.741 24.53 0.859 24.81 0.888 26.08 0.792 24.08 0.814 370.26 97.17
IDR [79] 2023 28.40 0.844 20.95 0.706 27.77 0.911 24.48 0.914 26.33 0.799 26.96 0.863 6.19 32.16
IRSDE [40] 2023 24.05 0.822 11.29 0.450 15.91 0.806 11.52 0.697 26.68 0.811 19.55 0.783 137.13 379.33
AutoDIR [20] 2024 29.32 0.863 15.65 0.707 15.31 0.706 19.01 0.829 28.47 0.864 22.43 0.799 115.86 63.38
DA-CLIP [41] 2024 28.63 0.854 19.50 0.730 28.23 0.934 27.26 0.941 26.47 0.818 27.54 0.881 32.96 158.14
GOUB [72] 2024 28.65 0.870 17.80 0.723 30.39 0.960 20.85 0.902 27.85 0.838 27.60 0.895 137.13 379.34
ConvIR [11] 2024 29.18 0.867 21.36 0.771 31.43 0.950 29.13 0.960 28.41 0.862 29.49 0.903 14.82 128.93
DeepSNNet [12] 2025 28.62 0.845 17.90 0.661 30.02 0.927 28.72 0.937 25.81 0.773 28.15 0.865 17.32 71.79
AWRaCLe [52] 2025 29.15 0.860 20.41 0.756 27.70 0.927 18.38 0.789 26.37 0.818 26.31 0.861 94.18 165.42
MaIR [30] 2025 29.45 0.864 21.76 0.750 30.80 0.955 30.39 0.960 28.28 0.859 29.51 0.904 20.71 110.44

### v2

@@ 1 Introduction
Universal image restoration emerges as a unified paradigm integrating the perception, representation, and elimination of diverse degradations [jiang2024survey, wang2025deep], with the aim of restoring high-quality images from the degraded low-quality observation. It typically encompasses a broad range of classical tasks, including denoising, deraining, dehazing, super-resolving, and others [chen2025towards, tian2024cross, kim2024frequency, wu2024seesr, goyal2024recent]. Owing to the high fidelity of restored details, it has been widely adopted as a precursor across various downstream tasks [marathe2022restorex, zhang2024ntire].
@@
Diffusion models have achieved remarkable advances in universal image restoration [luo2025taming]. Early methods [songdenoising, xia2023diffir, zhu2023denoising] follow the standard diffusion paradigm [ho2020denoising, karras2022elucidating] that maps images to a Gaussian distribution, and initializes the reverse inference from pure noise. Some approaches leverage generative priors pretrained from large models [luo2023diff, ma2025efficient] as conditional guidance for denoising networks. Others treat various restoration tasks as inverse problems by assuming access to degradation kernels [chungdiffusion, chung2023parallel, wang2025dgsolver, zhang2025improving]. However, the randomness of noise and reliance on specific priors compromise both stability and universality. Subsequent studies [luo2023image, liu2024residual] incorporate mean-reverting dynamics into diffusion stochastic differential equations (SDEs), clustering forward terminal states around degraded observations to retain task-relevant cues. Additionally, diffusion bridges [zhoudenoising] directly model point-to-point stochastic transitions between paired distributions, thereby strengthening data associations and improving restoration fidelity. Despite these advances, existing methods still rely on global noise perturbation to construct probabilistic trajectories, requiring rigid reverse denoising processes, as shown in Fig. 1. However, they fail to distinguish regions with varying degradation levels and imperfectly reconstruct intact regions, limiting restoration performance and adaptivity. Besides, a systematic and theoretical framework is absent to elucidate the intricate interconnections among existing diffusion bridge formulations.
In this work, we propose a scalable and unified diffusion bridge framework for universal image restoration, termed Residual Diffusion Bridge Model (RDBM), and conduct a comprehensive analysis of optimal distribution transitions. Specifically, we integrate the mean-reverting property of the Ornstein–Uhlenbeck SDEs with Doob’s $h$-transform [sarkka2019applied] to guide the terminal forward states toward degraded image distribution. Meanwhile, we use residuals from given distributions to dynamically modulate the probabilistic trajectories, thereby allowing the model to learn adaptive restoration of regions with varying degradation levels while mitigating redundant reconstruction in intact areas. Moreover, we theoretically demonstrate that our formulation yields the smooth distributional transition with respect to the residual-to-noise ratio. Building upon this, we uncover the mathematical essence of mainstream diffusion bridge formulations, all of which are special cases within our framework in specific configurations. Extensive experiments are conducted to verify the superiority of our method across diverse tasks including image restoration, translation, and inpainting both qualitatively and quantitatively. Our main contributions are summarized as follows:
@@ 2 Related Work
Denoising diffusion models [songscore, songdenoising] were initially developed for image generation. Methods such as DiffIR [xia2023diffir], DvSR [whang2022deblurring], and SR3 [zamir2020learning] directly repurpose diffusion models conditioned on degraded images for image restoration task, suffering from performance bottlenecks for task incompatibility. I2SB [liu20232] and ColdDiffusion [bansal2023cold] bypass explicit noise perturbations and instead learn a degraded diffusion process directly via the network. Besides, RDDM [liu2024residual], ReShift [yue2024efficient], ResFusion [shi2024resfusion], and DiffUIR [zheng2024selective] incorporate degradation distributions into the perturbation kernels to explicitly characterize degradation-aware diffusion processes. Moreover, IRSDE [luo2023image] employs a mean-reverting process to enforce diffusion trajectories that regress toward noisy degraded images with stationary variance. Additionally, DDBM [zhoudenoising], BBDM [li2023bbdm] and GOUB [yue2024image] further apply Doob’s $h$-transform to remove terminal noise, offering a tractable alternative to pave the probability path that connects degraded and clean images, thereby achieving remarkable restoration performance. Flow matching [lipman2023flow, liuflow] discards the stochastic noise and constructs the deterministic distribution transition path, thereby facilitating the optimal transport [zhu2025diffusion]. In contrast, our RDBM introduces the residual to modulate noise perturbation, enabling spatially adaptive restoration. Besides, RDBM can extend to these diffusion bridge models and flow matching in specific settings.
@@ 3.1 Diffusion Bridge Models
Diffusion SDEs [sohl2015deep, ho2020denoising] with drift term $\mathbf{f}(\cdot,t)$ and diffusion term $g(t)$ can be generally formulated as [karras2024analyzing]:
@@
where $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$ is score function. Furthermore, a diffusion process defined in Eq. (1) can be driven to arrive at a particular point of interest $\boldsymbol{\mu}$ via Doob’s $h$-transform [rogers2000diffusions]:
@@ 3.2 Ornstein Uhlenbeck Process
Figure 3: Overview of mainstream diffusion processes via SDEs, all of which are special cases of our framework. (a) OU process maps the data distribution to the prior distribution with noise. (b) OU bridge constructs probabilistic transition paths between given distributions. (c) Brownian bridge models linear expectations of intermediate states. (d) Our RDBM leverages residuals from paired distributions to adaptively modify the transition trajectories, maintaining a smooth residual-to-noise ratio.
@@ 4.1 Generalized Forward Process
which is governed by two terms. The first term depends on the residual component and fixed ratio $\lambda$. The second term is entirely determined by $\theta_{t}$ series and exhibits a monotonic decline, which diverges to infinity as time $t\!\rightarrow\!0$ and converges to an infinitesimal value as time $t\!\rightarrow\!T$. Previous works [yue2024image, zhoudenoising, luo2023image] typically set $\boldsymbol{\pi}=1$, thereby performing the global noise perturbation to uniformly disrupt the overall structure of images. This induces two ill-posed issues: (i) degraded regions with varying levels are treated equally and intact regions suffer redundant and imperfect reconstruction due to inevitable cumulative error in reverse process. (ii) pixel-wise numerator $x_{0}(i,j)\!-\!\boldsymbol{\mu}(i,j)]^{2}$ may exhibit discontinuous jumps, potentially distorting the smooth monotonic decay of the residual-to-noise ratio. Therefore, to maintain the dynamic equilibrium in transmission trajectories, we fix $\boldsymbol{\pi}=\mathbf{x}_{0}-\boldsymbol{\mu}$, thereby deriving our specific formulation within this framework with adaptive noise perturbation and pixel-independent residual-to-noise ratio:
@@ 4.2 Reverse Process and Training Objective
Assuming $p_{{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\boldsymbol{\mu})$ follows a Gaussian distribution centered at $m_{\dot{\theta}}(\mathbf{x}_{t},\mathbf{x}_{0})$ with a constant variance, minimizing the Kullback-Leibler divergence [kingmaauto] $D_{KL}$ is equivalent to reducing the distance between the means (see Suppl. G):
@@ 4.3 Analysis
We redefine a general mean-reverting process in Eq. (9) and employ Doob’s $h$ transform to derive the generalized diffusion bridge in Eq. (10) that exhibits the property of mean-arrival. Visualization comparisons of probability paths with several diffusion processes are illustrated in Fig. 3. We configure $\boldsymbol{\pi}$ to serve as the residual component for adaptive noise perturbation, yielding a smoothly decaying residual-to-noise ratio highly compatible with image restoration. Besides, other mainstream bridge models can be concluded in our framework, such as standard diffusion bridge [zhoudenoising], Brownian Bridge [li2023bbdm, liu20232], OU Bridge [yue2024image], Flow Matching [lipman2023flow, liuflow] and others, as summarized in Tab. 1. Please see Suppl. F for detailed derivations.
@@
$\theta_{t}\rightarrow 0$ $\lambda$ $\boldsymbol{\pi}=0$ Flow Matching [lipman2023flow, liuflow]
$\theta_{t}\rightarrow 0$ $\lambda\rightarrow\infty$ $\boldsymbol{\pi}=1$ VE Bridge [zhoudenoising]
$\theta_{t}\rightarrow 0$ $\lambda\rightarrow\frac{1}{2}$ $\boldsymbol{\pi}=1$ VP Bridge [zhoudenoising]
$\theta_{t}\rightarrow 0$ $2\lambda\theta_{t}\rightarrow 1$ $\boldsymbol{\pi}=1$ Brownian Bridge [li2023bbdm, liu20232]
$\theta_{t}$ $\lambda$ $\boldsymbol{\pi}=1$ OU Bridge [yue2024image]
@@ 7  Take the gradient descent step on
8 $\phantom{******}\nabla_{\theta}\|\boldsymbol{\pi}\epsilon-\boldsymbol{\pi}_{\epsilon}^{\dot{\theta}}(\mathbf{x}_{t},t,\boldsymbol{\mu})\|_{1}$
@@ 5.1 Datasets and Evaluation Metrics
Image deraining. We train our model on the merged datasets from Rain13K [Kui_2020_CVPR] and DeRaindrop [qian2018attentive], which cover diverse rain streaks and densities. Evaluation is conducted on both rain- and raindrop-removal tasks using the mixed datasets [yang2017deep, zhang2018density, qian2018attentive, li2022toward]. In addition, we assess zero-shot generalization on real-world datasets, including GT-Rain [ba2022gt-rain] without ground-truth for reference.
Low-light enhancement. We combine the LOL [wei2018deep] and VE-LOL-L [ll_benchmark] datasets, which furnish real and synthetic paired samples across diverse scenes with varying illumination and noise levels. Additionally, we employ the NPE [wang2013naturalness], MEF [7120119] and DICM [lee2013contrast] datasets to conduct zero-shot generalization on real-world scenarios.
Image desnowing. We adopt the CSD [chen2021all] dataset as the primary benchmark and evaluate real-world performance on Snow100K-Real [liu2018desnownet], which has no ground-truth.
Image dehazing. We adopt ITS_v2 [li2018benchmarking] and D-HAZY [Ancuti_D-Hazy_ICIP2016] as training benchmarks, encompassing diverse scenes under varying haze densities. The outdoor subset SOTS [li2018benchmarking] is used for evaluation, while real-world generalization is assessed on Dense-Haze [ancuti2019dense], NHRW [zhang2017fast], and NH-HAZE [NH-Haze_2020].
Image deblurring. We use the GoPro [nah2017deep] dataset to perform deblurring tasks, which contains various levels of blur obtained by averaging the clear images captured in very short intervals. To further validate the generalizability, we perform zero-shot restoration on the RealBlur [rim_2020_ECCV] dataset.
Benchmarks are evaluated using peak signal-to-noise ratio (PSNR) [huynh2008scope], structural similarity (SSIM) [wang2004image], natural image quality evaluator (NIQE) [lee2013contrast] in RGB space, and the learned perceptual image patch similarity (LPIPS) [snell2017learning] in feature space. For fairness, we compare our method with several universal restoration methods [zamir2022restormer, li2022all, potlapalli2023promptir, zhang2023ingredient, luo2023image, jiang2024autodir, luo2024controlling, yue2024image, cui2024revitalizing, deng2025deepsn, rajagopalan2025awracle, li2025mair, ma2023prores], which are all re-implemented on the mixed datasets for comparisons.
@@ 5.2 Implementation Details
Our method is trained using 8 Nvidia A800 GPUs with PyTorch [paszke2019pytorch] framework for 128h. Adam optimizer and L1 loss are employed for 500k iterations with a learning rate of $1\text{e}\!-\!4$. We set the batch size as $20$ and distribute it evenly to each task. We randomly crop patches of size $256\times 256$ from the original image as network input for training and use 10 timesteps for full-resolution testing. We utilize U-Net  [ronneberger2015u] architecture as network backbone. We change the channel number of the hidden layers $C$ to obtain different versions with varied parameter quantities:
@@ 5    $\mathbf{x}_{T}=\boldsymbol{\mu}$
Restomer [zamir2022restormer] 2022 28.54 0.847 21.75 0.742 28.53 0.919 26.54 0.924 26.44 0.799 27.61 0.869 26.09 140.99
AirNet [li2022all] 2022 24.78 0.774 13.05 0.485 25.80 0.885 18.53 0.827 25.76 0.782 24.01 0.809 5.76 301.27
Prompt-IR [potlapalli2023promptir] 2023 28.97 0.856 20.97 0.733 29.52 0.938 25.80 0.929 26.25 0.797 27.89 0.878 32.96 158.14
ProRes [ma2023prores] 2023 22.42 0.752 20.31 0.741 24.53 0.859 24.81 0.888 26.08 0.792 24.08 0.814 370.26 97.17
IDR [zhang2023ingredient] 2023 28.40 0.844 20.95 0.706 27.77 0.911 24.48 0.914 26.33 0.799 26.96 0.863 6.19 32.16
IRSDE [luo2023image] 2023 24.05 0.822 11.29 0.450 15.91 0.806 11.52 0.697 26.68 0.811 19.55 0.783 137.13 379.33
AutoDIR [jiang2024autodir] 2024 29.32 0.863 15.65 0.707 15.31 0.706 19.01 0.829 28.47 0.864 22.43 0.799 115.86 63.38
DA-CLIP [luo2024controlling] 2024 28.63 0.854 19.50 0.730 28.23 0.934 27.26 0.941 26.47 0.818 27.54 0.881 32.96 158.14
GOUB [yue2024image] 2024 28.65 0.870 17.80 0.723 30.39 0.960 20.85 0.902 27.85 0.838 27.60 0.895 137.13 379.34
ConvIR [cui2024revitalizing] 2024 29.18 0.867 21.36 0.771 31.43 0.950 29.13 0.960 28.41 0.862 29.49 0.903 14.82 128.93
DeepSNNet [deng2025deepsn] 2025 28.62 0.845 17.90 0.661 30.02 0.927 28.72 0.937 25.81 0.773 28.15 0.865 17.32 71.79
AWRaCLe [rajagopalan2025awracle] 2025 29.15 0.860 20.41 0.756 27.70 0.927 18.38 0.789 26.37 0.818 26.31 0.861 94.18 165.42
MaIR [li2025mair] 2025 29.45 0.864 21.76 0.750 30.80 0.955 30.39 0.960 28.28 0.859 29.51 0.904 20.71 110.44